# Andreas Petrou-Zeniou

### v3

@@ Appendix C Proofs of Other Theoretical Results
$\lim_{n\to\infty}Pr_{P}\left(\max_{\jmath\in\hat{J}_{R;n}}\frac{\left|\tilde{S}_{Y,\jmath}^{n}-\mu_{Y,\jmath}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}/n}}>\rho_{\alpha,\beta}(L^{n},U^{n},\hat{\Sigma}^{n}/n)\right)$
$\geq\lim_{n\to\infty}Pr_{P}\left(\max_{\jmath\in\hat{J}_{R;n}}\frac{\left|\tilde{S}_{Y,\jmath}^{n}-\mu_{Y,\jmath}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}/n}}>c_{1-\alpha}(J)+h\right)$ (40)
We know that, by the continuous mapping theorem, the term:
$\max_{\jmath\in\hat{J}_{R;n}}\frac{\left|\tilde{S}_{Y,\jmath}^{n}-\mu_{Y,\jmath\jmath}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}/n}}$
converges in distribution to:
$\max_{j\in J_{R}(P)}\frac{|\xi_{Y,jj}|}{\sqrt{\Sigma_{Y,jj}\left(P\right)}}$ (41)
where $\xi_{Y}$ is drawn from a multivariate Gaussian with mean zero and variance-covariance $\Sigma_{Y}(P)$, by the central limit theorem, continuous mapping theorem and the fact that the set $\hat{J}_{R;n}$ converges almost surely to $J_{R}(P)$. Thus, for any $h$:
$\lim_{n\to\infty}Pr_{P}\left(\max_{\jmath\in\hat{J}_{R;n}}\frac{\left|\tilde{S}_{Y,\jmath}^{n}-\mu_{Y,\jmath\jmath}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}/n}}>c_{1-\alpha}(J)+h\right)$
$=Pr_{P}\left(\max_{j\in J_{R}(P)}\frac{|\xi_{Y,jj}|}{\sqrt{\Sigma_{Y,jj}\left(P\right)}}>c_{1-\alpha}(J)+h\right)$ (42)
Now, since $h$ was arbitrary, it follows that:
@@
$\lim_{n\to\infty}Pr_{P}\left(\max_{\jmath\in\hat{J}_{R;n}}\frac{\left|\tilde{S}_{Y,\jmath}^{n}-\mu_{Y,\jmath\jmath}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}/n}}>\rho_{\alpha,\beta}(L^{n},U^{n},\hat{\Sigma}^{n}/n)\right)$
$\geq\lim_{n\to\infty}Pr_{P}\left(\max_{\jmath\in\hat{J}_{R;n}}\frac{\left|\tilde{S}_{Y,\jmath}^{n}-\mu_{Y,\jmath\jmath}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}/n}}>c_{1-\alpha}(J)\right)$ (43)
Ultimately implying (20) and thus proposition 5.3.
@@
Now, we provide proofs of propositions A.1 and A.2, which concern the validity of both locally simultaneous approaches to inference.
@@
Proof of Proposition A.1. Let $G$ be the event that $L\leq\Delta\leq U$. We know that $G$ satisfied $P_{\mu,\Sigma}(G)\geq 1-\beta_{2}$. On $G$, we have $p_{j}(L,U)\geq p_{j}$ for any $j=1,\ldots,p$. Consequently, we have that, on $G$, $J_{c}\subseteq\hat{J}_{c}$. We will also let $F$ be the even that $\hat{J}_{R}\subseteq J_{c}$, which satisfies $P_{\mu,\Sigma}(G)\geq 1-\beta_{1}$. Thus, we have:
$P_{\mu,\Sigma}\left(\exists\ \jmath\in\hat{J}_{R};\ \mu_{Y,\jmath}\not\in CS_{1-\alpha;\beta,\jmath}^{LS1}\right)$ $=P_{\mu,\Sigma}\left(\max_{\jmath\in\hat{J}_{R}}\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>c_{1-\alpha+\beta}\left(\hat{J}_{c}\right)\right)$
$\leq P_{\mu,\Sigma}\left(\max_{\jmath\in\hat{J}_{R}}\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>c_{1-\alpha+\beta}\left(\hat{J}_{c}\right)\cap G\cap F\right)+P_{\mu,\Sigma}(G^{c})+P_{\mu,\Sigma}(F^{c})$
$\leq P_{\mu,\Sigma}\left(\max_{\jmath\in\hat{J}_{R}}\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>c_{1-\alpha+\beta}\left(\hat{J}_{c}\right)\cap G\cap F\right)+\beta_{1}+\beta_{2}$
$\leq P_{\mu,\Sigma}\left(\max_{\jmath\in J_{c}}\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>c_{1-\alpha+\beta}\left(J_{c}\right)\cap G\cap F\right)+\beta_{1}+\beta_{2}$
$\leq P_{\mu,\Sigma}\left(\max_{\jmath\in J_{c}}\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>c_{1-\alpha+\beta}\left(J_{c}\right)\right)+\beta_{1}+\beta_{2}\leq\alpha-\beta+\beta_{1}+\beta_{2}=\alpha$
@@
Proof of Proposition A.2.
Let $J_{c}(\mu)$ be some set such that $P_{\mu,\Sigma}(\hat{J}_{R}\subseteq J_{c}(\mu))\geq 1-\beta$. We notice that $J_{c}(\mu)$ can be taken as follows:
$\left\{j;\ \min_{j^{\prime}\in J_{R}(\mu)}\frac{|\mu_{X,j}-\mu_{X,j^{\prime}}|}{\sqrt{\text{var}_{j,j^{\prime}}}}\leq d_{1-\beta}(\Sigma)\right\}$ (44)
Consequently, we notice that $J_{c}(\mu)$ depends on $\mu$ only through $J_{R}(\mu)$. By inverting 44, we can take a level $1-\beta$ confidence region for $J_{R}(\mu)$ as follows:
$\bar{J}_{R}(\mu):=\left\{j;\ \min_{j^{\prime}\in\hat{J}_{R}}\frac{|X_{j}-X_{j^{\prime}}|}{\sqrt{\text{var}_{j,j^{\prime}}}}\leq d_{1-\beta}(\Sigma)\right\}$ (45)
We notice that the $\hat{J}_{c}$ defined in 26 is simply the union of forecast intervals in 44 over all plausible $J_{R}(\mu)$ as collected in $\bar{J}_{R}(\mu)$. Thus, we have that $P_{\mu,\Sigma}(J_{c}\subseteq\hat{J}_{c})\geq 1-\beta$. Let $G$ be the event that $\max_{j,j^{\prime}\in R}\frac{|\xi_{X,j}-\xi_{X,j^{\prime}}|}{\sqrt{\text{var}_{j,j^{\prime}}}}\leq d_{1-\beta}(\Sigma)$. Moreover, we notice that the event $G$ implies that $J_{R}(\mu)\subseteq\bar{J}_{R}(\mu)$, that $J_{c}(\mu)\subseteq\hat{J}_{c}$ and that $\hat{J}_{R}\subseteq J_{c}(\mu)$. The first implication is a simple consequence of the triangle inequality. The second implication is a consequence of the first. For the final implication, we notice that $\max_{j,j^{\prime}\in R}\frac{|\xi_{X,j}-\xi_{X,j^{\prime}}|}{\sqrt{\text{var}_{j,j^{\prime}}}}\leq d_{1-\beta}(\Sigma)$ implies that:
$\forall\ j\in J_{R}(\mu),\ \exists\ j^{\prime}\in\hat{J}_{R}\quad\text{s.t.}\quad\frac{|X_{j}-X_{j^{\prime}}|}{\sqrt{\text{var}_{j,j^{\prime}}}}\leq d_{1-\beta}(\Sigma)$
$\implies\ \forall\ j\in\hat{J}_{R},\ \exists\ j^{\prime}\in J_{R}(\mu)\quad\text{s.t.}\quad\frac{|X_{j}-X_{j^{\prime}}|}{\sqrt{\text{var}_{j,j^{\prime}}}}\leq d_{1-\beta}(\Sigma)$
Indeed, we have on $G$ that $\hat{J}_{R}\subseteq J_{c}\subseteq\hat{J}_{c}$. Finally, we have:
$P_{\mu,\Sigma}\left(\exists\ \jmath\in\hat{J}_{R};\ \mu_{Y,\jmath}\not\in CS_{1-\alpha;\beta,\jmath}^{LS2}\right)$ $=P_{\mu,\Sigma}\left(\max_{\jmath\in\hat{J}_{R}}\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>c_{1-\alpha+\beta}\left(\hat{J}_{c}\right)\right)$
$\leq P_{\mu,\Sigma}\left(\max_{\jmath\in\hat{J}_{R}}\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>c_{1-\alpha+\beta}\left(\hat{J}_{c}\right)\cap G\right)+P_{\mu,\Sigma}(G^{c})$
$\leq P_{\mu,\Sigma}\left(\max_{\jmath\in J_{c}}\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>c_{1-\alpha+\beta}\left(J_{c}\right)\right)+P_{\mu,\Sigma}(G^{c})=\alpha-\beta+\beta=\alpha$
@@
Appendix D Inference on Multiple Winners in Microcredit
In this section, we present a stylized setting, based on the microcredit literature, wherein the inference on multiple winners problem becomes particularly relevant in assessing treatment effect heterogeneity across studies. We consider studies on microcredit treatment effects from Bosnia and Herzegovina (Augsburg et al. (2015)), Ethiopia (Tarozzi et al. (2015)), India (Banerjee et al. (2015a)), Mexico (Angelucci et al. (2015)), Mongolia (Attanasio et al. (2015)), and Morocco (Crépon et al. (2015)). All six studies are reviewed in Banerjee et al. (2015b). Depending on the outcome (either business profits, household earnings, or consumption), some of the above studies identify statistically significant treatment effects, while others do not. Pritchett and Sandefur (2015) suggest that this heterogeneity across studies is due to contextual differences between study sites. An alternative explanation is that those studies selected as statistically significant suffer from a winners’ curse, and that statistical significance arises due to a version of the multiple testing problem. We list the observational treatment effects of microcredit on household consumption from each of the six studies below:
Study Treatment Effect
Bosnia (Augsburg et al. (2015)) 0.193∗∗
Ethiopia (Tarozzi et al. (2015)) -0.0280
India (Banerjee et al. (2015a)) 0.0415
Mexico (Angelucci et al. (2015)) 0.0845∗∗
Mongolia (Attanasio et al. (2015)) -0.0317∗∗∗
Morocco (Crépon et al. (2015)) 0.194∗∗
Table 7: Observational Microcredit Treatment Effects on Household Income per Pritchett and Sandefur (2015)
We notice that, of the six treatment effects reported in Pritchett and Sandefur (2015), four are statistically significant at the 5% level, namely Augsburg et al. (2015), Angelucci et al. (2015), Attanasio et al. (2015), and Crépon et al. (2015). Empirically, we find that after correcting for cutoff-based selection, treatment effects remain statistically significant for all of these studies excluding Angelucci et al. (2015) and Attanasio et al. (2015). We demonstrate, via simulation, that a naive multiple testing correction may undercover following the selection described above, motivating a need for post-selection inference procedures. In these simulations, we compare three methods for inference: a naive bonferroni correction on only the selected programs, projection inference (as developed in Andrews et al. (2023), and our novel two-step inference procedure.
D.1 Empirical Results
We apply both our novel two-step confidence set and the naive projection confidence set to the four statistically significant studies identified in Pritchett and Sandefur (2015). In particular, we observe the following confidence intervals for the treatment effect in each selected study:
Study Treatment Effect Two-Step CI Projection CI Augsburg et al. (2015) 0.193 (0.005, 0.381) (0, 0.385) Crépon et al. (2015) 0.194 (0.006, 0.382) (0.002, 0.386) Angelucci et al. (2015) 0.084 (-0.103, 0.272) (-0.108, 0.277) Attanasio et al. (2015) -0.0317 (-0.219, 0.156) (-0.225, 0.161)
Table 8: Revised Observational Treatment Effects Confidence Sets for Statistically Significant Treatment Effects from Pritchett and Sandefur (2015)
Even after making the appropriate post-selection inference correction, we continue to observe statistically significant treatment effects in two out of the four selected studies, namely Augsburg et al. (2015) and Crépon et al. (2015). These findings suggest that, in part, the heterogeneity in treatment effects observed by Pritchett and Sandefur (2015) cannot be explained by the selection of statistically significant studies. Rather, treatment effect heterogeneity may indeed be due to variation in treatment design and setting.
D.2 Simulation Results
In this section, we develop a range of simulations based on the empirical setting introduced above. In particular, we consider three approaches to inference: projection accounting for selection, projection not accounting for selection, and the two-step approach. We provide further discussion of the first and third approaches in sections 4 and 5, respectively. The first approach is similar in spirit to a Bonferroni correction over all indices $J$. The second approach is similar in spirit to a Bonferroni correction over only the selected indices. Thus, the second approach can be thought of as correcting for the multiple testing component of the inference on multiple winners problem, but not the post-selection component. We find that projection accounting for selection provides wide marginal confidence intervals and overcovers, even asymptotically. On the other hand, projection not accounting for selection provides short marginal confidence intervals and undercovers in the finite sample, but provides valid coverage asymptotically. Finally, our novel two-step approach displays similar properties to projection accounting for selection, but does not overcover asymptotically. This is a key feature of the two-step approach, as discussed in proposition 5.3.
In the simulations that follow, we fix treatment effects normalized by standard errors, and select the studies for which this normalized value exceeds the critical value 1.96. We vary the variance covariance $\Sigma$ that results from the transformation described above, scaling by $1/n$ for a given sample size $n$. The figure below plots coverage probability for the three approaches described above with $n$ varying across the $x$-axis:
Figure 5: Coverage Probability in the Microcredit Application for Various Inference Procedures
We see reductions in over-coverage error by up to 50%. We see similar reductions in confidence set length. We consider the mean volume of our confidence sets across the three inference procedures described, finding that volume decreases with $n$, and that confidence set volume converges for the three methods asymptotically. In general, the two-step approach to inference provides shorter confidence sets than projection with a selection correction, while projection without a selection correction generally provides the smallest confidence sets in the finite sample.
Figure 6: Confidence Set Volume in the Microcredit Application for Various Inference Procedures
Appendix E Simulation Study Results

### v4

@@ Appendix C Proofs of Other Theoretical Results
$\lim_{n\to\infty}Pr_{P}\left(\max_{\hat{\jmath}\in\hat{J}_{R;n}}\frac{\left|Y_{\hat{\jmath}}-\mu_{Y,\hat{\jmath}}\right|}{\sqrt{\Sigma_{Y,jj}}}\leq\rho_{1-\alpha+\beta}(L^{n},U^{n};\hat{\Sigma}^{n})\right)\leq\lim_{n\to\infty}Pr_{P}\left(\max_{\hat{\jmath}\in\hat{J}_{R;n}}\frac{\left|Y_{\hat{\jmath}}-\mu_{Y,\hat{\jmath}}\right|}{\sqrt{\Sigma_{Y,jj}}}\leq c_{1-\alpha}(J)\right)$
Lemma s.6.1 of Romano and Shaikh (2012) and the continuous mapping theorem will give:
@@
$\lim_{n\to\infty}Pr_{P}\left(\max_{\hat{\jmath}\in\hat{J}_{R;n}}\frac{\left|Y_{\hat{\jmath}}-\mu_{Y,\hat{\jmath}}\right|}{\sqrt{\hat{\Sigma}^{n}_{Y,jj}}}\leq\rho_{1-\alpha+\beta}(L^{n},U^{n};\hat{\Sigma}^{n})\right)\leq\lim_{n\to\infty}Pr_{P}\left(\max_{\hat{\jmath}\in\hat{J}_{R;n}}\frac{\left|Y_{\hat{\jmath}}-\mu_{Y,\hat{\jmath}}\right|}{\sqrt{\hat{\Sigma}^{n}_{Y,jj}}}\leq c_{1-\alpha}(J)\right)$
Thus proving (20).
@@
Now, we provide proofs of propositions A.2 and 4.4, which concern the validity of both locally simultaneous approaches to inference.
@@
Proof of Proposition A.2. Let $G$ be the event that $L\leq\Delta\leq U$. We know that $G$ satisfied $P_{\mu,\Sigma}(G)\geq 1-\beta_{2}$. On $G$, we have $p_{j}(L,U)\geq p_{j}$ for any $j=1,\ldots,p$. Consequently, we have that, on $G$, $J_{c}\subseteq\hat{J}_{c}$. We will also let $F$ be the event that $\hat{J}_{R}\subseteq J_{c}$, which satisfies $P_{\mu,\Sigma}(G)\geq 1-\beta_{1}$. Thus, we have:
$P_{\mu,\Sigma}\left(\exists\ \hat{\jmath}\in\hat{J}_{R};\ \mu_{Y,\hat{\jmath}}\not\in CS_{1-\alpha;\beta,\hat{\jmath}}^{LSTS}\right)$ $=P_{\mu,\Sigma}\left(\max_{\hat{\jmath}\in\hat{J}_{R}}\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}>c_{1-\alpha+\beta}\left(\hat{J}_{c}\right)\right)$
$\leq P_{\mu,\Sigma}\left(\max_{\hat{\jmath}\in\hat{J}_{R}}\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}>c_{1-\alpha+\beta}\left(\hat{J}_{c}\right)\cap G\cap F\right)+P_{\mu,\Sigma}(G^{c})+P_{\mu,\Sigma}(F^{c})$
$\leq P_{\mu,\Sigma}\left(\max_{\hat{\jmath}\in\hat{J}_{R}}\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}>c_{1-\alpha+\beta}\left(\hat{J}_{c}\right)\cap G\cap F\right)+\beta_{1}+\beta_{2}$
$\leq P_{\mu,\Sigma}\left(\max_{\hat{\jmath}\in J_{c}}\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}>c_{1-\alpha+\beta}\left(J_{c}\right)\cap G\cap F\right)+\beta_{1}+\beta_{2}$
$\leq P_{\mu,\Sigma}\left(\max_{\hat{\jmath}\in J_{c}}\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}>c_{1-\alpha+\beta}\left(J_{c}\right)\right)+\beta_{1}+\beta_{2}\leq\alpha-\beta+\beta_{1}+\beta_{2}=\alpha$
@@
Proof of Proposition 4.4.
Our proof follows via application of theorem 1 of Zrnic and Fithian (2024b). We can define the following set of plausible targets, given a realization $X=x$:
$\hat{J}_{R}^{+}=\bigcup_{x^{\prime};\ \sup_{j^{\prime}\neq j}|(x_{j^{\prime}}^{\prime}-x_{j^{\prime}})-(x_{j}^{\prime}-x_{j})|\leq 2\bar{d}_{1-\beta}(\Sigma)}\hat{J}_{R}(x^{\prime})$
which we can verify is equivalent to $\hat{\Gamma}_{\beta}^{+}$ taking the acceptance region $A_{\beta}(\mu^{\prime})$ to be as follows:
$A_{\beta}(\mu^{\prime})=\left\{(y,x);\ \sup_{1\leq j,j^{\prime}\leq p}\left|X_{j}-\mu^{\prime}_{j}-(X_{j^{\prime}}-\mu_{j^{\prime}}^{\prime})\right|\leq\bar{d}_{1-\beta/2}(\Sigma)\right\}$
We may notice that $\sup_{j^{\prime}\neq j}|(x_{j^{\prime}}^{\prime}-x_{j^{\prime}})-(x_{j}^{\prime}-x_{j})|\leq 2\bar{d}_{1-\beta/2}(\Sigma)$ is equivalent to $\sup_{j^{\prime}\neq j}|x_{j^{\prime}}^{\prime}-x_{j}|\leq\bar{d}_{1-\beta/2}(\Sigma)$. Following the reasoning of theorem 3 in Zrnic and Fithian (2024b), we seek to find a most favorable choice of $y^{\prime}$ for an index $j$ to be selected. To allow $j\in\hat{J}_{R}(x^{\prime})$, let $j^{\prime}$ be the element in $\hat{J}_{R}(x)$ that minimizes $|x_{j}-x_{j^{\prime}}|$. Suppose, without loss, that $x_{j}<x_{j^{\prime}}$. We can obtain the most favorable perturbation by taking $x_{j}^{\prime}+\bar{d}_{1-\beta/2}(\Sigma)$ and $x_{j^{\prime}}^{\prime}-\bar{d}_{1-\beta/2}(\Sigma)$. Consequently, we can write:
$\hat{J}_{R}^{+}=\left\{j;\ \exists j^{\prime}\in\hat{J}_{R}\text{ s.t. }|X_{j^{\prime}}-X_{j}|\leq 2\bar{d}_{1-\beta/2}(\Sigma)\right\}$
@@
Finally, we provide a proof of proposition 4.5, which compares our two-step approach to inference to the locally simultaneous approach of Zrnic and Fithian (2024b).
Proof.
Proof of Proposition 4.5.
Notice that, whenever $\max_{j,j^{\prime}\in J}|\mu_{X,j}-\mu_{X,j^{\prime}}|\leq\bar{d}_{1-\beta}(\Sigma)$, the following series of implications holds, for arbitrary $j,j^{\prime}\in J$:
$|\xi_{X,j}-\xi_{X,j^{\prime}}|\leq\bar{d}_{1-\beta}(\Sigma)$
$\implies|\xi_{X,j}-\xi_{X,j^{\prime}}|+|\mu_{X,j}-\mu_{X,j^{\prime}}|\leq 2\bar{d}_{1-\beta}(\Sigma)$
$\implies|X_{j}-X_{j^{\prime}}|\leq 2\bar{d}_{1-\beta}(\Sigma)$
where the final inequality is a consequence of the triangle inequality. Consequently, $P_{\mu,\Sigma}(\max_{j,j^{\prime}\in J}|X_{j}-X_{j^{\prime}}|\leq 2\bar{d}_{1-\beta}(\Sigma))\geq 1-\beta$. On this event, it holds that $\hat{J}_{R}^{+}=J$. Moreover, on the event that $|\xi_{X,j}-\xi_{X,j^{\prime}}|\leq\bar{d}_{1-\beta}$, it holds that $L_{jj^{\prime}}\geq-3\bar{d}_{1-\beta}(\Sigma)$. Consequently, we notice that, with probability at least $1-\beta$, $\tilde{\rho}_{1-\alpha+\beta}(L)$ is bounded above by the $1-\alpha+\beta$ quantile of:
$\max_{j\in J}\frac{|\xi_{Y,j}|}{\sqrt{\Sigma_{Y,jj}}}\mathbb{1}\left(\xi_{X,j}-\xi_{X,j^{\prime}}\geq-3\bar{d}_{1-\beta}(\Sigma)\quad\text{for all}\quad j^{\prime}\in J\right)\leq\max_{j\in J}\frac{|\xi_{Y,j}|}{\sqrt{\Sigma_{Y,jj}}}$ (45)
where the inequality is strict with positive probability. It follows that $\tilde{\rho}_{1-\alpha+\beta}(L)<c_{1-\alpha+\beta}\left(\hat{J}_{R}^{+}\right)$ with probability at least $1-\beta$.
Appendix D Simulation Study Results